# Unified Approach to Power-Efficiency Trade-Off of Generic Thermal Machines

### v1

@@ Abstract
$P\leq\left(\frac{\rm{{I}_{\rm{rev}}}}{\mathcal{L}}\right)^{2}\frac{\eta\left(\eta_{\rm{rev}}-\eta\right)\left(1-\chi\eta_{\rm{rev}}\right)}{\theta\delta^{\frac{1+\theta}{2}}\chi^{-1}\left(1-\chi\eta\right)^{2}}\equiv P_{m}(\eta),$ (9)
where $\eta_{\rm{rev}}=\rm{O}_{\rm{rev}}/\rm{I}_{\rm{rev}}$ is the reversible efficiency. For a given efficiency, the power attains a maximum value when the cycle dissipation is minimized. By solving $dP_{m}(\eta)/d\eta=0$, the maximum power $P_{\rm{max}}$ and the corresponding EMP $\eta_{\rm{MP}}$ are obtained as
$\eta_{\rm{MP}}=\frac{\eta_{\rm{rev}}}{2-\chi\eta_{\rm{rev}}}$ (10)
@@
$P_{\rm{max}}=\theta\gamma\left(\frac{\rm{I}_{\rm{rev}}\eta_{\rm{rev}}}{2\mathcal{L}}\right)^{2}\delta^{-\frac{1+\theta}{2}}\left|\chi\right|.$ (11)
Normalize Eq. (9) with $P_{\rm{max}}$ and $\eta_{\rm{rev}}$, the main result of this Letter, Eq. (1), is obtained. We make two remarks here:
i) the $1/\tau$-scaling of dissipation can also be derived using the framework of linear irreversible thermodynamics [34, 9], hence our results are applicable to steady-state heat engines in liner response regime as well. Besides, utilizing the so-called shortcut to isothermality method [35, 32], the $1/\tau$-scaling can be maintained even beyond the slow-driving regime [33], making our results applicable to any thermal machine constructed using shortcut to isothermality and are not limited by driving speed. In the cases where the minimal cycle dissipation deviates from the $1/\tau$-scaling [36, 37, 38, 39, 40], for example, when $A_{\rm{min}}(\tau)\propto\tau^{\alpha}$ with $\alpha<0$ [36, 37, 38], Eq. (7) can directly yield analytical explicit PETOR. For more complex scenarios, with the series expansion of $A_{\rm{min}}$ respect to $\tau$ [39], our approach provides transcendental equation regarding $P$ and $\eta$, which will require further numerical calculation to obtain specific tradeoff relation.
iii) Equation (11) indicates that, for thermal machines with $\theta\gamma<0$ ($\theta=-\gamma=-1$ for heat pumps and $\theta=-\gamma=1$ for heat exchangers), the maximum power is negative, which is nonphysical. This implies that, under the condition of requiring positive output, the machines with $\theta\gamma<0$ do not have a locally optimal output power with respect to $\eta$, which results in a monotonic dependence of power on efficiency (PETORs with $\theta\gamma<0$ in Fig. 1). For thermal machines with $\theta\gamma>0$, the EMPs as functions of $\delta$ in the two-reservoirs case is plotted in Fig. 2(b). In this figure, $\tilde{\eta}_{\rm{MP}}=\eta_{\rm{MP}}/\eta_{\rm{rev}}\geq 0.5$ for heat engines while $\tilde{\eta}_{\rm{MP}}\leq 0.5$ for refrigerators, different curve styles correspond to different temperature ratios of reservoirs. In addition,Eq. (1) implies that once $\tilde{\eta}$ is fixed, the maximum $\tilde{P}$ is determined. In this scenario, increasing $P_{\rm{max}}$ can enhance P. As shown in Eq. (11), increasing $\rm{I}_{\rm{rev}}$ and decreasing $\mathcal{L}$ can effectively increase $P$, where the former can be achieved by enlarging the size of the machine (more working substance), and the latter can be realized through optimizing the driving path of the machine cycle [32, 41] or leveraging collective advantages offered by interactions inside working substances [42, 43].
PETORs for heat engines.— By specifying $\theta=1$ and $\gamma=1$, one has $\chi=\delta$ from Eq. (6), and thus, it follows from Eq. (1) that, the normalized PETOR for slow-driving heat engines reads
@@
When the scenario is further restricted to only two heat baths, a high-temperature one with temperature $T_{\rm{h}}$ and a low-temperature one with temperature $T_{\rm{c}}$, the reversible efficiency simplifies to its Carnot form, namely, $\eta_{\rm{rev}}=\eta_{\rm{C}}=1-T_{\rm{c}}/T_{\rm{h}}$. In this case, Eq. (13) is simplified as $\eta_{\rm{C}}/(2-\delta\eta_{\rm{C}})$ (see Fig. 2(b)), recovering the EMP of low-dissipation Carnot engine obtained by Esposito et al. [17]. Beside, by taking $\delta=1$, the overall upper bound for efficiency at arbitrary given power is obtained from Eq. (12) as (See Appendix B for details)
@@
and the overall lower bound, $\tilde{\eta}_{\rm{L}}=(1-\sqrt{1-\tilde{P}})/2$, is achieved with $\delta=0$. These bounds have been obtained before with different approaches [4, 6, 7] and experimentally tested [13]. Furthermore, in the case of two finite-sized heat reservoirs, the quasi-static efficiency of the heat engine is referred to as efficiency at maximum work ($\eta_{\rm{MW}}$[46, 47]. In this scenario, it can be easily verified that Eq.(12) is consistent with the result obtained with linear irreversible thermodynamics, as given by [Eq. (10)] of Ref. [9]. The EMP of the engine satisfies $\eta_{\rm{MW}}/2\leq\eta_{\rm{MP}}\leq\eta_{\rm{MW}}/(2-\eta_{\rm{MW}})$ [34, 9]. We do not delve into detailed discussions on this issue here.
PETORs for refrigerators.— For refrigerators, their efficiency is referred to as the Coefficient of Performance (COP), typically denoted by the symbol $\varepsilon$. It follows from Eq. (1) that, by specifying $\theta=-1$ and $\gamma=-1$, i.e., $\chi=-(1-\delta)^{-1}$, and using variable substitution $\eta\rightarrow\varepsilon$ and $\eta_{\rm{rev}}\rightarrow\varepsilon_{\rm{rev}}$, the PETOR for slow-driving refrigerators is obtained as
@@
where $\varepsilon_{\rm{rev}}\equiv Q_{\rm{rev}}^{\rm{c}}/W_{\rm{rev}}$ represents the COP a refrigerator operating reversibly. In this situation, the efficiency at maximum cooling power and corresponding maximum cooling power follows
@@
respectively. Since $(1-\delta)^{-1}\in[1,\infty)$, the EMP of refrigerators satisfy $0\leq\varepsilon_{\rm{MP}}\leq\varepsilon_{\rm{rev}}/(2+\varepsilon_{\rm{rev}})$. For given reversible cycle qualities, the maximum cooling power and its corresponding COP are solely determined by the dissipation ratio $\delta$. We note that as the dissipation at the high-temperature end diminishes ($\delta\rightarrow 1$), the cooling power diverges, whereas the COP approaches zero at this point. However, the product of the two is a constant, namely, $\lim_{\delta\rightarrow 1}\left(P_{\rm{max}}\epsilon_{\rm{MP}}\right)=[W_{\rm{rev}}\varepsilon_{\rm{rev}}/(2\mathcal{L})]^{2}$, indicating that when the dissipation at the high-temperature heat reservoir is extremely small, the optimized extreme point moves along a hyperbola. In more specific case with two heat baths, one has $\eta_{\rm{rev}}=\epsilon_{\rm{C}}=T_{\rm{c}}/\left(T_{\rm{h}}-T_{\rm{c}}\right)$ and corresponding EMP is $\epsilon_{\rm{MP}}=\epsilon_{\rm{C}}/[2+(1-\delta)^{-1}\epsilon_{\rm{C}}]$, which is consisting with Ref. [22]. In Fig. 2(b), $\epsilon_{\rm{MP}}$ as a function of $\delta$ with $T_{\rm{c}}/T_{\rm{h}}=0.1,0.5,0.8$ is plotted as the curves below $\tilde{\eta}_{\rm{MP}}\leq 0.5$, larger $\epsilon_{\rm{C}}$ (larger $T_{\rm{c}}/T_{\rm{h}}$) leads to higher $\epsilon_{\rm{MP}}$.
PETORs for heat pumps and heat exchangers.— From Eq. (9), the PETOR of heat pump is (see Fig. 1)
@@
As we mentioned in remark iii) below Eq. (11), heat pumps and heat exchangers with $\theta\gamma<0$ do not have locally maximum positive output power. The monotonicity of their PETORs can be directly observed from Fig. 1. Mathematically, $\theta\gamma<0$ leads to $P$ being a monotonically increasing function of cycle time $\tau$ in the physically meaningful region of $\tau>0$, without any optimal time that locally maximizes $P$. This is also revealed in the specific analysis of heat pumps in Ref. [23].
It is worth noting that, in real world circumstance, practical heat exchangers usually do not have complete thermodynamic cycles; instead, they achieve rapid heat transfer by allowing the high-temperature substance in the heat exchanger to come into contact with the low-temperature environment to be heated though some medium [48, 49]. With thermodynamic cycles, the output mechanical work will reduce the heat release to the cold reservoir, thereby lowering the efficiency of the heat exchangers. However, local heat exchange around the heat exchanger may also lead to temperature non-uniformity in a relatively large space to be heated. It can be envisioned that by converting some heat into mechanical work, it is possible to perform work on the medium in the heated environment such as air, accelerating heat convection and achieving temperature uniformity. This analysis suggests that designing heat exchangers with adjustable mechanical output power (essentially a type of heat engine) to balance heat transfer efficiency and temperature uniformity in the heated environment is a practical task worth exploring, which has implications for guiding the optimal design of heat exchangers from the perspectives of engineering thermodynamics.
Summary.—We have obtained power-efficiency trade-off relation applicable to generic thermal machines. The core of our approach lies in two key points: i) unifying the performance description of different thermal machines using sign functions, and ii) establishing relations between total cycle dissipation, cycle time, and machine’s power and efficiency, the trade off of which results from the lower bound on dissipation qualified by the thermodynamic length. Our results are applicable to cases with multiple heat baths as well as finite-sized heat reservoirs. Specifically, in the case of two heat reservoirs, by selecting appropriate cycle parameters, the presented general PETOR naturally recovers typical PETORs such as those for low-dissipation Carnot heat engines and Carnot refrigerators, and that for irreversible heat engines with finite-sized reservoirs. The EMP determined by our PETOR also recovers several existing results obtained case by case. Moreover, our work is highly generalizable (see remarks below Eq. (11)), Eq. (1) is applicable to any thermal machine driven by isothermal shortcut protocols [33], and the approach we use to derive PETOR can also be extended to study thermodynamic cycles that do not satisfy the $1/\tau$-scaling of dissipation [36, 37, 38, 39, 40]. This study lays the foundation for a unified description of thermodynamic cycles and general approaches to thermal machine performance optimization.

### v2

@@ III.1 Trade-off relations between power and efficiency
$P\leq\left(\frac{\mathrm{I}_{\mathrm{rev}}^{\alpha+1}\chi}{\Sigma\theta\delta^{\frac{1+\theta}{2}}}\right)^{\frac{1}{\alpha}}\frac{\eta\,(\eta_{\mathrm{rev}}-\eta)^{\frac{1}{\alpha}}\,(1-\chi\eta_{\mathrm{rev}})}{(1-\chi\eta)^{\frac{\alpha+1}{\alpha}}}\equiv P_{m}(\eta),$ (8)
where $\eta_{\rm{rev}}=\rm{O}_{\rm{rev}}/\rm{I}_{\rm{rev}}$ is the reversible efficiency. The equal sign in this unified power-efficiency trade-off relation is saturated when the cycle dissipation is minimized. Figure 2 intuitively demonstrate the generality of Eq. (8) across different thermal machines. This example ($\alpha=1$) will be discussed in detail in Sec. IV.1.
Figure 2: The normalized power-efficiency trade-off relation ($\tilde{\eta}\equiv\eta/\eta_{\mathrm{rev}},~\tilde{P}\equiv P/P_{\mathrm{max}}$) of four thermal machines. The blue dashed (red dash-dotted) curve represents the Eq. (8) with $\delta\rightarrow 1$($\delta\rightarrow 0$), the normalized reversible efficiency $\tilde{\eta}=1$ is plotted with the green dotted line, and the shadow area marks the available operation region of the machines. In this plot, $\alpha=1$ and $\eta_{\rm{rev}}=0.5$ are used.
Furthermore, it follows from Eq. (8) that the normalized power-efficiency trade-off relations for heat engines ($\theta=\gamma=1$) are illustrated in Fig. 3 for two limiting cases of $\delta=0,1$. As shown in this figure, when $\alpha$ increases, the extreme point on the right-hand side of the trade-off relation gradually approaches $(1,1)$, indicating that the heat engines can approach reversible efficiency at maximum power. This can be regarded as a thermodynamic advantage [43]. We will further discuss the properties of the EMP in Sec. III.2. Notably, the dissipation $A\propto\tau^{-\alpha}$ can be achieved by smoothly modulating the system-reservoir interaction, where any $\alpha>0$ can be attained through appropriate modifications of the interaction [44]. Particularly, $\alpha=1$ corresponds to a standard isothermal process with constant system-reservoir interaction [9, 44, 15]. The parameter $\alpha$ is introduced not only to generalize the expression of our unified approach, but also to provide an additional degree of freedom for optimizing the performance of thermal machines. This degree of freedom offers a more general perspective for enhancing thermodynamic processes and enables broader control over the operational characteristics of such systems.
Figure 3: The normalized trade-off relations of heat engines associated with $\delta\to 0$ [(a)] and $\delta\to 1$ [(b)] with different $\alpha$ and $\eta_{\rm{rev}}=0.5$. The blue dotted, red dash-dotted, and purple dashed curve represent $\alpha=1,4,50$, respectively, $\tilde{\eta}=1$ is plotted with the green dotted line.
By specifying parameter $\theta=\gamma=-1$, we demonstrate the trade-off relations of refrigerators in Fig. 4 with different $\delta$ and $\alpha$. As $\delta$ increases, the maximum power point on the right-hand side of the trade-off relation moves downward and gradually approaches (1,0). This differs from the $\delta$-monotonicity observed in heat engines, owing to the distinct definition of efficiency. Moreover, when $\delta\nrightarrow 1$, increasing $\alpha$ causes the entire trade-off relation to shift upward, similar to the behavior of heat engines. Specifically, in the Fig. 4(d), for large $\alpha$ and $\delta=1$, the trade-off forms a trapezoid–a behavior not previously observed. This indicates that the selection of EMP appears to be arbitrary. In the $\eta-P$ diagram, the long-time regime occupies the upper section of the curve, while the short-time regime resides in the lower section. A critical time threshold exists within the short-time regime: below this threshold, power increases sharply, but refrigerating performance remains low and exhibits strong time dependence. Above the threshold, power flattens as irreversibility diminishes significantly with increasing duration, driving substantial efficiency enhancements.
Figure 4: The normalized trade-off relations associated with $\alpha=1$[(a)], $\alpha=2$[(b)], $\alpha=5$[(c)], and $\alpha=50$[(d)]. The blue dotted, red dash-dotted, and purple dashed curves are plotted with $\delta=0,0.7,1$, respectively. In this plot, $\eta_{\rm{rev}}=0.5$.
After examining the derivative formula of $P_{m}(\eta)$ given in Eq. (8), it was observed that when the efficiency approaches zero or reaches the reversible efficiency $\eta_{\mathrm{rev}}$, both the initial value of the derivative and its rate of change become significantly large. This phenomenon causes the trade-off relation curve between $P$ and $\eta$ to rapidly decline to zero near these two extreme points. In contrast, in the middle region of the efficiency regime, the derivative remains relatively small and nearly constant, resulting in minimal variation in $P$ with respect to $\eta$. Consequently, a trapezoid pattern emerges in the graphical representation.
III.2 Efficiency at maximum power
Figure 5: Normalized EMPs of heat engines ($\tilde{\eta}_{\rm{MP}}\equiv\eta_{\mathrm{MP}}/\eta_{\mathrm{rev}}$) as functions of $\tilde{\eta}_{\rm{rev}}$. (a) The blue dotted, red dash-dotted, yellow solid, and purple dashed curves are plotted with $\alpha=1,2,10,100$, respectively, and $\delta=1$ is fixed. (b) The blue dotted, red dash-dotted, and purple dashed curves are associated with $\delta=0,1/(1+\sqrt{1-\eta_{\mathrm{rev}}}),1$, respectively, and $\alpha=1$ is fixed.
As the extreme point on the power-efficiency trade-off relation [9], EMP [23, 22, 24, 25, 13, 20] is a crucial parameter for evaluating finite-time heat engines ($\theta=\gamma=1$) and refrigerators ($\theta=\gamma=-1$). It follows from Eq. (8) that by solving $dP_{m}(\eta)/d\eta=0$, the corresponding EMP is
$\eta_{\rm{MP}}=\frac{\alpha\eta_{\rm{rev}}}{1+\alpha-\chi\eta_{\rm{rev}}}.$ (9)
Combining Eqs. (9) and (5), the Normalized EMP $\tilde{\eta}_{\rm{MP}}\equiv\eta_{\mathrm{MP}}/\eta_{\mathrm{rev}}$ of heat engines as a function of $\eta_{\mathrm{rev}}$ is plotted in Fig. 5 with different $\alpha$ [(a)] and $\delta$ [(b)]. It is shown in this figure that, for given $\eta_{\mathrm{rev}}$, increasing either $\alpha$ or $\delta$ can improve the EMP. Besides, it is worth mentioning that, due to the different signs (positive or negative) of $\chi$ corresponding to different thermal machines, the monotonicity of their EMP with respect to $\delta$ also varies, see Fig. 6 for a special case ($\alpha=1$) of refrigerators [(a)] and heat engines[(b)].
Figure 6: Normalized EMPs of refrigerators [(a)] and heat engines [(b)] as functions of $\delta$ in the case of $\alpha=1$. The blue dotted curve, red dash-dotted curve, and purple dashed curve are plotted with $\eta_{\rm{rev}}=0.2,0.5,0.9$, respectively. The green solid line represents $\tilde{\eta}=0.5$.
Furthermore, the leading-term coefficient of EMP with respect to $\eta_{\mathrm{rev}}$ is a noteworthy issue in the optimization of finite-time heat engines [24, 23]. In the near equilibrium regime, this coefficient is $1/2$ [24, 16], irrespective of engine details. However, such universality hinges on the validity of linear irreversible thermodynamics, which may be violated away from equilibrium [43]. Here, as a result of Eq. (9),
$\eta_{\rm{MP}}(\eta_{\rm{rev}}\ll 1)\approx\frac{\alpha}{1+\alpha}\eta_{\rm{rev}},$ (10)
which means that in our setup, the $1/2$-universality holds exclusively when $\alpha=1$. This fact is explicitly demonstrated in Fig. 6(b): for heat engines with different $\delta$ under $\alpha=1$, their EMP converges exactly to $\eta_{\rm{MP}}=\eta_{\mathrm{rev}}/2$ when $\eta_{\mathrm{rev}}\to 0$. In contrast, when $\alpha\to\infty$, EMP approaching $\eta_{\rm{rev}}$ follows directly from Eq.(10), and this trend is clearly shown in Fig. 5(a).
IV Specific trade-off relations for different $\alpha$
As a specific application of the general results obtained above, we focus on two typical scenarios in this section: the slow-driving isothermal regime ($\alpha=1$) and the quantum finite-time adiabatic regime ($\alpha=2$), and derive the specific forms of trade-off relations for several thermal machines.
IV.1 The slowing driving isothermal regime of $\alpha=1$
In the slow-driving finite-time isothermal processes, the minimal energetic dissipation is proved to be $A_{\rm{min}}=\mathcal{L}^{2}\tau^{-1}$ [14], where $\mathcal{L}$ is the thermodynamic length in thermodynamic geometry theory. This lower bound for irreversibility is achieved when the driving protocol of the cycle maintains uniform speed along the thermodynamic length trajectory (in the parametric space defined by thermodynamic state variables) [47, 14, 52, 54]. With this $1/\tau$-scaling (i.e., $\alpha=1$), Eq. (8) becomes
$P\leq\left(\frac{\rm{{I}_{\rm{rev}}}}{\mathcal{L}}\right)^{2}\frac{\eta\left(\eta_{\rm{rev}}-\eta\right)\left(1-\chi\eta_{\rm{rev}}\right)}{\theta\delta^{\frac{1+\theta}{2}}\chi^{-1}\left(1-\chi\eta\right)^{2}},$ (11)
the corresponding EMP $\eta^{\alpha=1}_{\rm{MP}}$ and maximum power $P^{\alpha=1}_{\rm{max}}$ are
$\eta^{\alpha=1}_{\rm{MP}}=\frac{\eta_{\rm{rev}}}{2-\chi\eta_{\rm{rev}}}$ (12)
@@ IV.1 The slowing driving isothermal regime of $\alpha=1$
$P^{\alpha=1}_{\rm{max}}=\theta\gamma\left(\frac{\rm{I}_{\rm{rev}}\eta_{\rm{rev}}}{2\mathcal{L}}\right)^{2}\delta^{-\frac{1+\theta}{2}}\left|\chi\right|,$ (13)
respectively. Therefore, the normalized trade-off relation for a generic thermal machine operating in the slow-driving regime is
where $\tilde{P}\equiv P/P^{\alpha=1}_{\rm{max}}$ and $\tilde{\eta}\equiv\eta/\eta_{\rm{rev}}$. We make two remarks here:
i) the $1/\tau$-scaling of dissipation can also be derived using the framework of linear irreversible thermodynamics [38, 55, 11], hence our results are applicable to steady-state heat engines in linear response regime as well. Besides, utilizing the so-called shortcut to isothermality method [56, 52], the $1/\tau$-scaling can be maintained even beyond the slow-driving regime [54], making our results applicable to any thermal machine constructed using shortcut to isothermality and not limited by driving speed.
ii) As a result of Eq. (13), the maximum power of thermal machines with $\theta\gamma<0$ ($\theta=-\gamma=-1$ for heat pumps and $\theta=-\gamma=1$ for heat exchangers) is negative, which is nonphysical. This implies that, under the condition of requiring positive output, the machines with $\theta\gamma<0$ do not have a locally optimal output power with respect to $\eta$, which results in a monotonic dependence of power on efficiency (trade-off relations with $\theta\gamma<0$ in Fig. 2). For thermal machines with $\theta\gamma>0$, the EMPs as functions of $\delta$ in the two-reservoirs case are plotted in Fig. 6. In this figure, $\tilde{\eta}_{\rm{MP}}=\eta_{\rm{MP}}/\eta_{\rm{rev}}\geq 0.5$ for heat engines while $\tilde{\eta}_{\rm{MP}}\leq 0.5$ for refrigerators, different curve styles correspond to different $\eta_{\rm{rev}}$. In addition, Eq. (14) implies that once $\tilde{\eta}$ is fixed, the maximum $\tilde{P}$ is determined. In this scenario, increasing $P_{\rm{max}}$ can enhance $P$. As shown in Eq. (13), increasing $\rm{I}_{\rm{rev}}$ and decreasing $\mathcal{L}$ can effectively increase $P$, where the former can be achieved by enlarging the size of the machine (more working substance), and the latter can be realized through optimizing the driving path of the machine cycle [52, 18] or leveraging collective advantages offered by interactions inside working substances [57, 43].
It follows from Eq. (11) that several typical trade-off relations for various thermal machines can be directly obtained, as demonstrated in Fig. 2. By specifying $\theta=1$ and $\gamma=1$, one has $\chi=\delta$ from Eq. (5). Thus, it follows from Eq. (14) that, the normalized trade-off relation for slow-driving heat engines reads
@@
When the scenario is further restricted to only two heat reservoirs, a high-temperature one with temperature $T_{\rm{h}}$ and a low-temperature one with temperature $T_{\rm{c}}$, the reversible efficiency simplifies to its Carnot form, namely, $\eta_{\rm{rev}}=\eta_{\rm{C}}=1-T_{\rm{c}}/T_{\rm{h}}$. In this case, Eq. (16) is simplified as $\eta_{\rm{C}}/(2-\delta\eta_{\rm{C}})$, recovering the EMP of low-dissipation Carnot engine obtained by Esposito et al. [25]. The EMP in this case has been plotted in Fig. 5(b), where the blue dotted and purple dashed curves represent the upper bound $\eta_{\rm{C}}/(2-\eta_{\rm{C}})$ ($\delta=1$) and lower bound $\eta_{\rm{C}}/2$ ($\delta=0$), respectively. The red dash-dotted is associated with $\delta=1/(1+\sqrt{1-\eta_{\mathrm{C}}})$, representing the Curzon-Ahlborn efficiency [22, 13]
Beside, by taking $\delta=1$, the overall upper bound for efficiency at arbitrary given power is obtained from Eq. (15) as (See Appendix B for details)
@@
and the overall lower bound, $\tilde{\eta}_{\rm{L}}=(1-\sqrt{1-\tilde{P}})/2$, is achieved with $\delta=0$. These bounds have been obtained before with different approaches [6, 8, 9] and experimentally tested [16]. Furthermore, in the case of two finite-sized heat reservoirs, the quasi-static efficiency of the heat engine is referred to as efficiency at maximum work ($\eta_{\rm{MW}}$[60, 61]. In this scenario, it can be easily verified that Eq. (15) is consistent with the result obtained with linear irreversible thermodynamics, as given by [Eq. (10)] of Ref. [11]. The EMP of the engine satisfies $\eta_{\rm{MW}}/2\leq\eta_{\rm{MP}}\leq\eta_{\rm{MW}}/(2-\eta_{\rm{MW}})$ [46, 55, 11]. We do not delve into detailed discussions on this issue here.
For refrigerators, their efficiency is referred to as the COP, denoted by the symbol $\varepsilon$. It follows from Eq. (14) that, by specifying $\theta=-1$ and $\gamma=-1$, i.e., $\chi=-(1-\delta)^{-1}$, and using variable substitutions $\eta\rightarrow\varepsilon$ and $\eta_{\rm{rev}}\rightarrow\varepsilon_{\rm{rev}}$, the trade-off relation for slow-driving refrigerators is obtained as
@@
where $\varepsilon_{\rm{rev}}\equiv Q_{\rm{rev}}^{\rm{c}}/W_{\rm{rev}}$ represents the COP a refrigerator operating reversibly. In this situation, the efficiency at maximum cooling power and the corresponding maximum cooling power follow
@@
respectively. Since $(1-\delta)^{-1}\in[1,\infty)$, the EMP of refrigerators satisfies $0\leq\varepsilon_{\rm{MP}}\leq\varepsilon_{\rm{rev}}/(2+\varepsilon_{\rm{rev}})$. For given reversible cycle qualities, the maximum cooling power and its corresponding COP are solely determined by the dissipation ratio $\delta$. We note that as the dissipation at the high-temperature end diminishes ($\delta\rightarrow 1$), the cooling power diverges, whereas the COP approaches zero at this point. However, the product of the two is a constant, namely, $\lim_{\delta\rightarrow 1}\left(P_{\rm{max}}\epsilon_{\rm{MP}}\right)=[W_{\rm{rev}}\varepsilon_{\rm{rev}}/(2\mathcal{L})]^{2}$, indicating that when the dissipation at the high-temperature heat reservoir is extremely small, the optimized extreme point moves along a hyperbola. In more specific case with two heat baths, one has $\eta_{\rm{rev}}=\epsilon_{\rm{C}}=T_{\rm{c}}/\left(T_{\rm{h}}-T_{\rm{c}}\right)$ and the corresponding EMP is $\epsilon_{\rm{MP}}=\epsilon_{\rm{C}}/[2+(1-\delta)^{-1}\epsilon_{\rm{C}}]$, which is consisting with Ref. [30]. In Fig. 6(a), $\epsilon_{\rm{MP}}$ as a function of $\delta$ with $\eta_{\rm{C}}=0.2,0.5,0.9$ is plotted as the curves below $\tilde{\eta}_{\rm{MP}}\leq 0.5$, larger $\epsilon_{\rm{C}}$ leads to lower $\tilde{\epsilon}_{\rm{MP}}$.
From Eq. (11), the trade-off relation of heat pump is (see Fig. 2)
@@
As we mentioned in remark ii) below Eq. (14), heat pumps and heat exchangers with $\theta\gamma<0$ do not have locally maximum positive output power. The monotonicity of their trade-off relations can be directly observed from Fig. 2. Mathematically, $\theta\gamma<0$ leads to $P$ being a monotonically increasing function of cycle time $\tau$ in the physically meaningful region of $\tau>0$, without any optimal time that locally maximizes $P$. This is also revealed in the specific analysis of heat pumps in Ref. [31].
IV.2 The finite-time quantum adiabatic regime of $\alpha=2$
It has been proved that $\alpha=2$ corresponds to the finite-time adiabatic processes where $A_{\rm{min}}=\Sigma/\tau^{2}$ [41]. In this case, Eq. (8) becomes
$P\leq\frac{\rm{{I}^{3/2}_{\rm{rev}}}\eta\left(1-\chi\eta_{\rm{rev}}\right)}{\sqrt{\Sigma\theta\delta^{1+\theta}}}\sqrt{\frac{\chi\left(\eta_{\rm{rev}}-\eta\right)}{\left(1-\chi\eta\right)^{{}^{3}}}}.$ (21)
This serves as the trade-off relation for finite-time quantum Otto thermal machines, which has not been obtained before, to the best of our knowledge.
Figure 7: The normalized trade-off relations associated with $\alpha=2$ for heat engines[(a-b)] and refrigerators[(c-d)]. The blue dotted, red dash-dotted, and purple dashed curve are plotted with $\eta_{\mathrm{rev}}=0.2,0.6,0.9$, respectively. $\tilde{\eta}=1$ is plotted with the green dotted line. $\delta=0.1$ is used to plot (a) and (c), while $\delta=1$ for (b) and (d).
Figure 7 illustrates trade-off relations with different $\eta_{\mathrm{rev}}$ and $\delta$. As can be seen from Fig. 7(a-b), the differences in trade-off relations of heat engines caused by different $\eta_{\mathrm{rev}}$ increase with the increase of $\delta$. Meanwhile, a larger $\delta$ leads to a higher EMP of heat engines. In contrast, increasing $\delta$ shifts the trade-off relations of refrigerators downward and reduces their EMPs, as shown in Fig. 7(c-d). This implies that heat engines and refrigerators require different optimization strategies with respect to $\delta$. In addition, substituting $\alpha=2$ into Eq. (9) yields
$\eta_{\rm{MP}}=\frac{2\eta_{\rm{rev}}}{3-\chi\eta_{\rm{rev}}},$ (22)
which is consistent with the result of finite-time quantum Otto heat engines ($\chi=\delta$) given in Ref. [41]. By comparing this result with Eq. (16), the fact that the quantum Otto engine can exceed the Carnot cycle’s EMP with the same $\delta$ indicates that out-of-equilibrium quantum thermodynamic cycles may have greater advantages in finite-time optimization under specific conditions [41, 51, 42].
V Conclusion
In this work, we have proposed a unified framework for describing arbitrary thermal machines, addressing a key limitation in previous studies: the requirement for separate performance parameter definitions for machines with distinct functions, which caused practical inconvenience. The core of our approach lies in two key points: i) unifying performance descriptions across different thermal machines using sign functions, and ii) establishing relations among total cycle dissipation, cycle time, and the machine’s power and efficiency. This unified framework is theoretically applicable to any form of finite-time machine dissipation $A(\tau)$.
The typical dissipation $A(\tau)\propto\tau^{-\alpha}$ draws critical benefits from control strategies, namely optimizing the regulation scheme of the system-environment interaction [44]. Leveraging the unified thermal machines description, we derived a general power-efficiency trade-off relation [Eq.(8)] for generic thermal machines exhibiting $\tau^{-\alpha}$-type cycle irreversibility. We further demonstrate the controllability of this trade-off relation in different cases with specified parameters $\alpha$, $\theta$, and $\gamma$. Our general result recovers several previous results [Sec. IV], yields the complete performance curve, and enables the investigation of a broader range of operating regimes for thermal machines. Notably, when $\alpha$ approaches infinity, theoretical derivations indicate that thermal machines can simultaneously attain reversible efficiency and maximum power[Sec. III.2]. This outcome not only validates the universal analytical capability of this method for thermal machine behavior across different scaling regimes but also makes innovative predictions regarding the thermodynamic characteristics under extreme scaling limits. We have also determined the trade-off relation of finite-time quantum adiabatic Otto engines[Sec. IV.2] for the first time. This study lays the foundation for a unified description of thermodynamic cycles and general approaches to thermal machine performance optimization.
As a final remark, the right-angled triangle trade-off relation shown in Fig. 3(a) is similar to the optimal trade-off relation for the degenerate heat engine [43]. This suggests that regulating the thermal machine-reservoir coupling (current work and Ref. [44]) and regulating the intrinsic structure of the machine’s working substance (Ref. [43]) may have a more profound connection in the context of optimizing machine performance.
Acknowledgment